Find the exponential equation for a graph containing these points (0,-2), (1,3) and (-5,-3)
Answer by ankor@dixie-net.com(22740)

Find the exponential equation for a graph containing these points (0,-2), (1,3) and (-5,-3)
:
Using the form a^x + b = y
0, -2
a^0 + b = -2
1 + b = -2
b = -2 - 1
b = -3
:
1, -3
a^1 - 3 = 3
a = 3 + 3
a = 6
:
Our equation: y = 6^x - 3
:
See if the works with -5,-3
y = 6^-5 -3
y = .0001286 - 3, (6^-5 on a graph is not distinguishable from 0)
y ~ -3